# Towards Global AI Inclusivity: A Large-Scale Multilingual Terminology Dataset (GIST)

### v5

@@ A Large-Scale Multilingual Terminology Dataset (GIST)
Jiarui Liu,1thanks: Equal contribution. Iman Ouzzani,1footnotemark: Wenkai Li,1footnotemark: Lechen Zhang,2 Tianyue Ou,1 Houda Bouamor,1 Zhijing Jin,3 Mona Diab1 1Carnegie Mellon University, 2University of Michigan, 3University of Toronto {jiaruil5, iouzzani, wenkail}@andrew.cmu.edu, zjin@cs.toronto.edu, mdiab@andrew.cmu.edu
@@ 6.2 Experiment Setup
The second evaluation set is manually created by combining text from two sources: 50 held-out AI research papers and 50 model cards generated by Liu et al. (2024). From this set, we randomly sampled 500 English text chunks from it for evaluation. To create ground truth labels for this set, we use Google Translate, a state-of-the-art machine translation model (Zhu et al., 2023; Santosa et al., 2024), to produce initial translations. We then prompt GPT-4o to refine these translations by updating AI terminology. This evaluation set is designed to explore the application of our AI terminology dictionary in two major domains: AI research papers and model cards. Both evaluation set references have been verified by human experts and exhibit high quality. Further details are provided in Section C.4.
@@ 6.3 Experiment Results
The word alignment method demonstrates mixed performance, showing improvements for Chinese and Japanese translations but leading to declines for Arabic, French, and Russian. This discrepancy is due to linguistic differences: Chinese exhibits minimal morphological changes, allowing straightforward substitution of terminology with limited disruption to surrounding syntax; In contrast, languages like Arabic often require agreement in gender, number, and syntactic roles, making noun replacement more complex and error-prone. These findings underscore the superior performance of the prompt-powered refinement method, as well as the importance of tailoring terminology integration approaches to the linguistic characteristics of target languages. See Sections D.1 and D.2 for further discussion of the results. We further validate our findings using a one-sided paired t-test in Section D.3.
@@ B.4 Dataset Quality Assessment Details
To evaluate the quality of translations in our dataset, we conducted two tasks involving human annotators. Annotators were presented with pairs of translations generated by different methods and were tasked with evaluating their relative quality. The analysis of their ratings is summarized in Tables 10, 11, and 12. Below, we describe the findings in detail.

### v6

@@ A Large-Scale Multilingual Terminology Dataset (GIST)
Jiarui Liu,1thanks: Equal contribution. Iman Ouzzani,1footnotemark: Wenkai Li,1footnotemark: Lechen Zhang,2 Tianyue Ou,1 Houda Bouamor,1 Zhijing Jin,3,4,5 Mona Diab1 1Carnegie Mellon University, 2University of Michigan, 3University of Toronto 4Max Planck Institute for Intelligent Systems, Tuebingen, Germany 5Vector Institute {jiaruil5, iouzzani, wenkail}@andrew.cmu.edu, zjin@cs.toronto.edu, mdiab@andrew.cmu.edu
@@ 6.2 Experiment Setup
The second evaluation set is manually created by combining text from two sources: 50 held-out AI research papers and 50 model cards generated by Liu et al. (2024). From this set, we randomly sampled 500 English text chunks from it for evaluation. To create ground truth labels for this set, we use Google Translate, a state-of-the-art machine translation model (Zhu et al., 2023; Santosa et al., 2024), to produce initial translations. We then prompt GPT-4o to refine these translations by updating AI terminology. This evaluation set is designed to explore the application of our AI terminology dictionary in two major domains: AI research papers and model cards.
To validate the quality of our evaluation set references, we conducted a manual evaluation involving five expert annotators who assessed 100 randomly sampled examples (50 from each evaluation set). Each annotator was presented with the source English sentence, its translation, and the relevant AI terminologies and their translations. They completed three tasks: (1) rating overall translation quality on a scale from 1 (nonsensical) to 5 (excellent), (2) assessing grammatical and morphological correctness (Yes/No), and (3) evaluating the accuracy of terminology translation (Yes/No). As summarized in Table 13, results indicate high annotation consistency and translation quality across all five languages. Average translation scores ranged from 3.11 (Russian) to 3.96 (Chinese), and binary accuracy for grammar and terminology typically exceeded 80%. These findings demonstrate the robustness and linguistic reliability of our evaluation sets and affirm the quality of GPT-4o-assisted refinement. Further details are provided in Section C.4.
@@ 6.3 Experiment Results
The word alignment method demonstrates mixed performance, showing improvements for Chinese and Japanese translations but leading to declines for Arabic, French, and Russian. This discrepancy is due to linguistic differences: Chinese exhibits minimal morphological changes, allowing straightforward substitution of terminology with limited disruption to surrounding syntax; In contrast, languages like Arabic often require agreement in gender, number, and syntactic roles, making noun replacement more complex and error-prone. These findings underscore the superior performance of the prompt-powered refinement method, as well as the importance of tailoring terminology integration approaches to the linguistic characteristics of target languages. See Sections D.1 and D.2 for further discussion of the results.
We further validate our findings using a one-sided paired t-test in Section D.3, which confirms that the prompt-powered refinement method significantly outperforms both direct translation and the word alignment approach across all evaluated languages.
@@ 8 Conclusion
In addition, expanding the terminology dataset automatically presents unique challenges. The rapidly evolving nature of the AI field requires frequent updates to ensure the GIST dataset remains current and comprehensive. At present, identifying new, domain-relevant terms still depends heavily on human expertise to assess whether a term is both novel and impactful within the AI research community. While large language models can aid in extracting terminology and generating candidate terms, their utility is constrained by fixed knowledge cutoffs, limiting their ability to capture the most recent developments. As a result, fully automating the update process remains a complex and unresolved task.
@@ Acknowledgments
This material is based in part upon work supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by Schmidt Sciences SAFE-AI Grant; by NSERC Discovery Grant RGPIN-2025-06491; by a National Science Foundation award (#2306372); by a Swiss National Science Foundation award (#201009) and a Responsible AI grant by the Haslerstiftung.
@@ B.4 Dataset Quality Assessment Details
To evaluate the quality of translations in our dataset, we conducted two tasks involving human annotators. Annotators were presented with pairs of translations generated by different methods and were tasked with evaluating their relative quality. The analysis of their ratings is summarized in Tables 2, 3, and 12. Below, we describe the findings in detail.